OpenStudy (ericwise):

How to factor the trinomial: x^2 - x - 14 ?

OpenStudy (ericwise):

I'm lost at the part where you need to find the corresponding product and sum values

OpenStudy (ericwise):

None of the factors seem to work out

OpenStudy (jhannybean):

Use the completing the square method,

Nnesha (nnesha):

find two number if you multiply them you should get -14 and if you add or subtract them you should get -1 this is headphone method

OpenStudy (ericwise):

Hmm, I've forgotten what that one's all about @Jhannybean

OpenStudy (ericwise):

@Nnesha, yeah I considered that. But it seems that none of the factors 1, 2, 7, 14 have a sum or difference of -1

OpenStudy (ericwise):

well, except 1 and 2... doesnt work though for the product

OpenStudy (ericwise):

wow, all this attention... hi everyone! :D

OpenStudy (anonymous):

Did we teach you this the other day?

OpenStudy (ericwise):

nope, must've been another one @wio, I don't think I've addressed a completing the square factoring question before

OpenStudy (jhannybean):

That's one method to go about factoring, but i'm sure there are several others.

OpenStudy (jhannybean):

Do you understand what I did, @EricWise ?

OpenStudy (ericwise):

I do not seem to recognize what is happening in the second step

Nnesha (nnesha):

i guess something wrong :(

OpenStudy (anonymous):

Where did \(-1/2\) come from?

OpenStudy (jhannybean):

I think I've doubled it. Whoops.

OpenStudy (jhannybean):

Sorry, I misread the function.

OpenStudy (jhannybean):

Now in the completing the square method, you want to "complete" a quadratic function inside the parenthesis.

OpenStudy (ericwise):

We're assuming that the 16 is actually a 14 right? :)

OpenStudy (jhannybean):

Yes, lol

OpenStudy (jhannybean):

\[x^2-x-14\]

OpenStudy (jhannybean):

\[(x^2-x)-14\]

OpenStudy (jhannybean):

In order to do that, we find a new \(c\) value, because if you compare the quadratic formula to what is inside the parenthesis, you will see that you have: \(ax^2+bx\) whereas in order to complete a quadratic, you must fit the form :\(ax^2+bx+c\)

OpenStudy (jhannybean):

Therefore we're essentially completing a quadratic within the parenthesis.

OpenStudy (jhannybean):

\[c=\left(\frac{b}{2}\right)^2\]

OpenStudy (jhannybean):

Our \(b\) is \(-1\). \[c=\left(\frac{-1}{2}\right)^2 = \frac{1}{4}\]

OpenStudy (jhannybean):

With this \(c\) value, we have completed our quadratic. rewrite the function. \[\left(x^2-x+\frac{1}{4}\right)-14-\frac{1}{4}\]

OpenStudy (ericwise):

oh ok... this might be ringing a bell

OpenStudy (anonymous):

\[ x^2 - x - 14 \\ ax^2+bx+c \]

OpenStudy (jhannybean):

now to keep our equation consistent with our original, we must also subtract a \(-1/4\)

OpenStudy (jhannybean):

Like what @wio posted above my post.

OpenStudy (jhannybean):

A little trick to remember is: \[c=\color{red}{\left(\frac{-1}{2}\right)}^2 = \frac{1}{4}\] Our quadratic will always simplify to \(x\pm \text{the red portion}\)

OpenStudy (ericwise):

oh okay!

OpenStudy (jhannybean):

\[\left(x-\frac{1}{2}\right)^2 -\frac{57}{4}\]

OpenStudy (jhannybean):

that is your factored form.

OpenStudy (jhannybean):

What I did earlier, was I mistook your equation for \(x^2-\dfrac{x}{2} -14\)

OpenStudy (ericwise):

oh okay^^

OpenStudy (jhannybean):

Do you understand the completing the square method? :)

OpenStudy (ericwise):

So if this whole factored thing is all equal to zero, that means: